# itemizestditemize

### v1

@@ VII Evaluation in Road Detection and Coverage
We evaluate RAG in two scenarios of simulated experiments of road detection and coverage tasks, one involving 15 and the other 45 robots. In both cases, the robots operate fully as a distributed/mesh network. RAG achieves planning at the order of $1$Hz and has superior coverage performance against the competitive near-optimal algorithms.
To perform the evaluations, we extend AirSim [4] to the multi-robot setting, simulating v2v communication delays (Fig. 3). We use ROS1. Thus, the evaluations include ROS1’s inherent time delays due to the transport layer and message passing overhead, as reported in Figs. 710.
The code of the simulator will be made available herein.
Common Simulation Setup across Simulated Scenarios. We first define the task of road detection and coverage (Fig. 1), then introduce the compared algorithms, and, finally, present the simulation pipeline (Fig. 3).
@@
Multiple aerial robots with onboard cameras are deployed in an unknown urban environment and tasked to detect and monitor the roads (Fig. 1). To perform the task, given the currently visible environment, the robots jointly plan how to move per the collaborative autonomy pipeline in Fig. 2. Particularly, the task takes the form of the optimization problem in eq. 1 where $f$ denotes the number of road pixels captured by all robots’ collective FOV after they traverse their planned trajectories $\{a_{i}\}_{i\,\in\,{\cal N}}$$f$ is non-decreasing, submodular, and 2nd-order submodular [39]— and ${\cal V}_{i}$ denotes robot $i$’s available trajectories at the current planning step. Specifically, ${\cal V}_{i}$ defines the possible directions that the robot can move in, and the speed the robot can move with. We assume that every robot can move in any of the 8 cardinal directions —N E S W NE SE NW SW— relative to its body frame, for $10$ m at $3$ m/s.
Without loss of generality, the deployed robots are assumed to have the same onboard sensing and communication capabilities: all robots are equipped with (i) an inertial measurement unit (IMU); (ii) a GPS signal receiver; (iii) a downward-facing camera mounted on a gimbal that enables the camera to point to any of the 8 cardinal directions relative to the robot’s body frame; and (iv) a communication module that enables $10$ MByte/s data rate per inter-robot communication channel. Each robot can establish a few communication channels per a specified communication bandwidth constraint, only with those robots that are within $100$ m.
@@
Across various bandwidth constraints for the robots, we compare RAG with two competitive near-optimal algorithms: the Sequential Greedy (SG) algorithm [18], also known as Coordinate Descent [1], and its state-of-the-art Depth-First-Search variant (DFS-SG[21]. Each algorithm is tested in 30 trials, each lasting 2 minutes. In more detail, the setup is as follows:
We test RAG for different bandwidth constraints that vary from 00 up to $6$. In each case, the same bandwidth constraint applies to all robots. Each such version of RAG is denoted by RAG-$k$nn, where $k=0,\dots,6$. For each $k$, the communication network is determined by having each robot select $k$ nearest other robots as neighbors subject to the $100$ m communication range. If fewer than $k$ others are within the communication range, then all are selected as neighbors.
@@
The DFS-SG algorithm enables SG to be applied to networks that are not necessarily a line graph, but the networks still need to be strongly connected. To this end, we construct strongly connected graphs by first randomly constructing line graphs as for SG, then randomly adding a few undirected edges to the line graphs, particularly, 30 edges for the 15-robot case and 90 edges for the 45-robot case. At each planning round, DFS-SG randomly picks the first robot to select an action, and the order of all other robots is determined by a distributed method based on depth-first search. The resulting decision sequence may involve relay robots that transmit information between robots that are not directly connected and, thus, DFS-SG generally requires longer decision times than SG.
@@
  • Network Self-Configuration: This module applies only to RAG since only RAG enables the network to self-configure itself across the planning steps subject to the robots’ bandwidth constraints and the robots’ relative locations. Particularly, as described also in the above paragraph for RAG-$k$nn, at the beginning of each planning step, each robot selects its $k$ nearest other robots within its communication range. This neighborhood selection scheme is justified by Example 1.
  • Perception: This process is different for each algorithm since each algorithm requires each robot to process information received from different sets of robots. For RAG, the robots need the following three operations to detect the nearby environment and evaluate where would be best to move: (i) At the beginning of each planning step, each robot hovers at a constant height (30 m) and rotates its camera to take a picture in each of the 8 cardinal directions relative to its body frame. The captured FOV in each direction is of size $34.6\times 26.0$ m2. Then, the robot uses semantic segmentation to detect the road segments in each of the 8 captured images. The resulting segmented image has size $1$ MByte. (ii) Each robot, upon receiving segmented images and relative poses from neighbors that have committed actions in specific directions, stitches these images together. That way, the robot reconstructs the collective FOV of its neighbors that have committed an action. This reconstruction will be used next by the robot to select in which direction to move to cover the most new road area. To this end, (iii) the robot stitches its 8 captured images respectively with the previous reconstruction and counts the extra road pixels that can be covered in each of the corresponding 8 cardinal directions.
@@
  • Action Coordination: This process is different for each algorithm. For RAG, given their neighbors’ already committed actions, the robots that are still deciding will first each pick a trajectory whose FOV gives the most amount of newly covered road area. Then, they will compare this amount with one another. Those who win their neighbors will commit to their picked trajectories and share the corresponding FOV and camera pose with neighbors. Otherwise, they will receive the FOVs and camera poses selected by newly committed robots and repeat the process above.
@@
Figure 7: Total Road-Area Coverage vs. Time (15 robots). The total road area covered is normalized by the result of RAG-6nn. The reported time includes ROS1 delays, with a maximum of 1.971 seconds per replanning period.
Figure 8: Maximum Total Road-Area Coverage vs. Replanning Time (15 robots). The total road area covered is normalized by the result of RAG-6nn. Team replanning times include ROS1 delays, with a maximum of 1.971 seconds per replanning period.
@@ VII-A Evaluations with 15 Robots
We present the results for the scenario with 15 robots. To this end, we first present in more detail the simulation setup.
Simulation Setup. Across the Monte Carlo tests, the robots are deployed near to one another in a small area such as for the leftmost group in Fig. 1, such that their FOVs largely overlap.
@@
Results. The results are summarized in Figs. 7 and 8. We observe that RAG achieves real-time planning and superior coverage. We also observe that a smaller neighborhood in favor of higher decision efficiency results in only a diminishing impact on the coverage performance. These observations validate the theoretical observation in Remarks 7 and 8.
Coverage performance over time
In Fig. 7, we present the total road area covered across time for each of the compared algorithms. We observe that RAG-$k$nn achieves in all but one case of $k$ ($k=0$) superior performance, one to two orders faster. For example, RAG-$1$nn achieves the same performance at $t=40$ sec as SG and DFS-SG do at $t=140$ sec. The lower final performance for $k=0$ is expected since then there is no coordination. Nevertheless, for the first $10$ to $20$ sec, RAG-00nn achieves better coverage since no coordination enables higher replanning frequency.
The reason that the algorithms SG and DFS-SG underperform is that they require more time to terminate per planning step, therefore, over the allocated $120$ sec, the robots replan fewer times in comparison to RAG.
Coverage vs. replanning time
In Fig. 8, we plot the total road area covered by the end of the simulation period ($130$ sec) vs. the average replanning time. We observe that when the neighborhood size for RAG increases, the covered road area increases, as the replanning time also does. This is expected since larger neighborhoods imply more coordination, thus better coverage performance but slower pace.
Smaller neighborhoods can be enough
In Fig. 8, we observe diminishing returns in coverage performance as the neighborhood size increases: although the total road area covered increases significantly from $k$ to $k+1$ for $k=0,1,2,4$, it increases marginally for higher $k$. Therefore, centralizing coordination beyond a size of neighborhoods $k=4$ appears inefficient: RAG-4nn seems to balance the trade-off of total coverage and decision speed, enjoying both fast replanning ($3.3$ sec) and high coverage rate (95%): RAG-4nn achieves 95% coverage at $1/3$ of the replanning time of RAG-6nn.
The observation validates the theory that decision speed increases proportionally with the size of the network (Theorem 3) but the total utility sublinearly (Proposition 2).
VII-B Experiments with 45 Drones
We demonstrate the scalability of our algorithm from 15 to 45 robots. We show that RAG maintains similar decision time and coverage as for the 15-robot setting, being at most 2 times slower, due to enabling parallelized decision-making. In contrast, the compared algorithms now require $2$ to $3$ minutes per replanning step instead of $12$ seconds, a $15$ times increase. By executing the experiments, we also demonstrate that the provided simulator pipeline can support such large-scale multi-robot scenarios.
Simulation Setup. Across the Monte Carlo tests, the robots are divided into three groups, with the robots within each group being deployed near to one another such that their FOVs largely overlap (Fig. 1). In such a setting, RAG automatically enables parallelized decision-making, maintaining similar decision time and coverage as for the 15-robot setting, in contrast to the state-of-the-art near-optimal algorithms.
@@ VII-B Experiments with 45 Drones
Results. The results are summarized in Figs. 9 and 10. We observe that RAG still maintains real-time planning and superior coverage, with all qualitative observations from the 15-robot setting applying also here. The near-optimal algorithms now require on average $2$ to $3$ minutes per replanning step, a $15$ times increase compared to the 15-robot case, as expected due to their superlinear time complexity (Table I).
Figure 9: Total Road-Area Coverage vs. Time (45 robots). The total road area covered is normalized by the result of RAG-6nn. The reported time includes ROS1 delays, with a maximum of 5.808 seconds per replanning period.
Figure 10: Maximum Total Road-Area Coverage vs. Replanning Time (45 robots). The total road area covered is normalized by the result of RAG-6nn. Team replanning times include ROS1 delays, with a maximum of 5.808 seconds per replanning period.
VIII Conclusion and Future Work
Summary. We provided a rigorous coordination algorithm that enables teams of distributed mobile robots to self-configure their communication topology to achieve real-time and near-optimal coordination. Our coordination paradigm is in contrast to the current paradigms that are either near-optimal but impractical for replanning times or real-time but offer no near-optimality guarantees. We made theoretical and algorithmic contributions to characterize and balance this trade-off between decision speed and optimality. On the theoretical side, we provided an analysis of how the network topology at the robot level —each robot’s coordination neighborhood— affects the near-optimality of the coordination at the global level. On the algorithmic side, we provided a communication- and computation-efficient algorithm that enables the agents to balance the trade-off. Our algorithm is up to two orders faster than competitive near-optimal algorithms. In realistic simulations of surveillance tasks with up to 45 robots, the algorithm enabled real-time planning at the order of $1$ Hz and superior coverage performance. To enable realistic simulations, we provided a high-fidelity simulator that extends the AirSim simulator to integrate a collaborative autonomy pipeline and simulate v2v communication delays.
@@ VIII Conclusion and Future Work
Future Work. We will enhance our results in three directions. (i) RAG assumes synchronous communication. Although RAG can be trivially modified to handle such cases, e.g., by instructing each robot to execute its action without first waiting for all other robots to select actions, its near-optimality guarantees provided in this paper may become invalid. Our future work will extend our theoretical and algorithmic analysis beyond the above limitations. (ii) We will also extend our results to handle effective task execution over long time horizons. For example, in collaborative mapping over long time horizons, the team needs to stay updated on the areas that have been mapped, such that the current plans do not repeat past actions. This is in addition to the focus of this paper that only the current plans among the robots do not overlap. To this end, we may need also to handle network connectivity constraints over the time horizon. (iii) Finally, we will enhance our simulator by integrating the simulation of realistic communication channels and protocols towards communication-aware and -efficient coordination algorithms.
@@
$=f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}\cup{\cal B}_{2}})-f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}})$ (15)
$\leq f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}})-f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}})=0,$ (16)
where the inequality holds since $f$ is submodular. Hence, ${\sf coin}_{f,i}({\cal N}_{i})$ is non-increasing in ${\cal N}_{i}$. Therefore, for any $i\in{\cal N}$, ${\sf coin}_{f,i}({\cal N}_{i})$ achieves the lower bound ${\sf coin}_{f,i}({\cal N}_{i})\geq{\sf coin}_{f,i}({\cal N}\setminus\{i\})=f(a_{i})-f(a_{i}\,|\,\emptyset)\geq 0$. For the upper bound,
@@
$\leq f({\cal A}^{\sf OPT}\cup{\cal A}^{{\sf RAG}})$ (19)
$=f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf OPT}\,|\,{\cal A}^{{\sf RAG}}\cup{\cal A}^{\sf OPT}_{[i-1]})$ (20)
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf OPT}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (21)
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (22)
$=2f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$
$\hskip 142.26378pt-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$ (23)
$=2f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$
$\quad\quad\quad\quad\quad\quad\quad-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]}\cup{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}})$ (24)
$\leq 2f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}})$ (25)
$\leq 2f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{{\cal N}_{i}^{c}})$ (26)
$\leq 2f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,{\sf coin}_{f,i}({\cal N}_{i}),$ (27)
where eq. 19 holds from the monotonicity of $f$; eqs. 20 and 23 result from telescoping the sums; eqs. 21 and 26 hold from the submodularity of $f$; eq. 22 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily; eq. 24 holds since ${\cal A}^{{\sf RAG}}_{[i-1]}={\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]}\cup{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}}$; eq. 25 holds from the 2nd-order submodularity of $f$; and eq. 27 holds from Definition 3. Therefore, eq. 9 holds.
Then, in the fully centralized case where ${\cal N}_{i}^{c}=\emptyset,\forall i\in{\cal N}$,
$f({\cal A}^{\sf OPT})$ $\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})$ (28)
$\leq f({\cal A}^{{\sf RAG}})+\sum_{i\,\in\,{\cal N}}\,f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})$ (29)
$=2f({\cal A}^{{\sf RAG}}),$ (30)
where eq. 28 follows eq. 22, eq. 29 holds since ${\cal N}_{i}={\cal N}\setminus\{i\},\forall i\in{\cal N}$, and eq. 30 holds by telescoping the sums. Thus, eq. 10 is proved.
Alternatively, we can prove eq. 10 with eq. 27 and using ${\sf coin}_{f,i}({\cal N}_{i})=f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG})=0$.
@@
where eq. 32 holds from eq. 5, and eq. 33 holds from [51, Lemma 2.1]. Combining eqs. 27 and 33, the lower bound in eq. 12 can be proved. ∎

### v2

@@ VII Evaluation in Road Detection and Coverage
We evaluate RAG in the provided simulator across four scenarios of road detection and coverage. The scenarios span two team sizes (15 and 45 robots), and two communication rates (0.25 Mbps), and 100 Mbps) —up to three orders of magnitude faster than competitive near-optimal submodular optimization algorithms— and, when the robots maintain a neighborhood size $\geq 2$, superior mean coverage. The experiments also demonstrate RAG scales linearly with the network size, as predicted by our theoretical analysis: from the 15-robot to the 45-robot case, RAG scales linearly, being at most 3x slower compared to the 15-robot case. In contrast, the compared near-optimal algorithms scale cubically, being up to 60x slower compared to the 15-robot case.
Common Simulation Setup across Simulated Scenarios. We first present the task of road detection (Fig. 1), then the compared algorithms and the simulation pipeline (Fig. 3).
@@
Multiple aerial robots with onboard cameras are deployed in an unknown urban environment and tasked to maximize the total new road area detected after each action coordination step (Fig. 1). The environment is unknown to the robots (no map is available a priori), necessitating the robots to use their onboard cameras to perform exploration for road detection.
The robots are deployed close to each other relative to the size of their FOVs and, as a result, their FOVs may often overlap (Fig. 1). Therefore, coordination becomes necessary for the robots to spread in the environment such that they detect different road segments and maximize the total new road area detected after each action coordination step.
To perform the task, given the currently visible environment, the robots coordinate how to move per the collaborative autonomy pipeline in Fig. 2. Particularly, the task takes the form of the optimization problem in 1 where $f$ denotes the number of road pixels captured by all robots’ collective FOV after they traverse their agreed trajectories $\{a_{i}\}_{i\,\in\,{\cal N}}$$f$ is non-decreasing, submodular, and 2nd-order submodular [44]— and ${\cal V}_{i}$ denotes robot $i$’s available trajectories at the current coordination round. ${\cal V}_{i}$ defines the possible directions that the robot can move in, and the speed the robot can move with. For simplicity, we assume that every robot can move in any of the 8 cardinal directions —N E S W NE SE NW SW— relative to its body frame, for $10$ m at $3$ m/s.
Without loss of generality, the deployed robots are assumed to have the same onboard sensing and communication capabilities: All robots are equipped with (i) an inertial measurement unit (IMU); (ii) a GPS signal receiver; (iii) a downward-facing camera mounted on a gimbal that enables the camera to point to any of the 8 cardinal directions relative to the robot’s body frame; and (iv) a communication module —either a Digi XBee 3 Zigbee 3.0 ($0.25$ Mbps) or Silvus Tech SL5200 ($100$ Mbps)— for inter-robot communication. Each robot can establish a few communication channels with robots within $100$ m range, subject to bandwidth constraints.
Figure 7: Comparative Analysis of Coverage Performance with 15 robots for $100$ Mbps data rate. The reported times include ROS1 delays of up to 0.18 sec per action coordination step. Particularly, in Figs. 710, all coverage data are normalized by the mean results of the corresponding RAG-7nn instances, and the shared area and cross-hairs represent 1 standard deviation for y-axis in (a) and both axes in (b).
Figure 8: Comparative Analysis of Coverage Performance with 15 robots for $0.25$ Mbps data rate. The reported times include ROS1 delays of up to 0.81 sec per action coordination step.
@@
Across various bandwidth constraints for the robots, we compare RAG with two competitive near-optimal algorithms: the Sequential Greedy (SG) algorithm [11], also known as Coordinate Descent [1], and its state-of-the-art Depth-First-Search variant (DFS-SG[31]. The algorithms are commonly used for information gathering with multiple robots (e.g., see the papers [1, 15, 17] and the survey [5]. In more detail, the setup is as follows:
We test RAG for different bandwidth constraints varying from 00 up to $7$. [^8] Although neighborhood sizes larger than 7 are possible in the simulator, we limit them to at most size 7 to achieve a reasonable duration for completing each Monte Carlo trial. Particularly, the neighborhoods’ scalability to $k>7$ is constrained by the msgpack-rpc protocol implementation over TCP/IP via rpclib. The current use of a single TCP port for multi-drone API calls creates a communication bottleneck, increasing client response times. For example, our experiments with 45 robots, which include 30 trials that simulate the robots’ evolution for 500 sec for 9 different algorithms (RAG $k$-nn for $k=1,\ldots,7$, SG, and DFS-SG), require 3 days for each algorithm to be tested. In each case, the same bandwidth constraint applies to all robots. Each such version of RAG is denoted by RAG-$k$nn, where $k=0,\dots,7$. For each $k$, the communication network is heuristically determined by having each robot select $k$ nearest other robots as neighbors subject to the $100$ m communication range. If fewer than $k$ others are within the communication range, then all are selected as neighbors.
@@
The DFS-SG algorithm enables SG to be applied to networks that are not necessarily a line graph, but the networks still need to be strongly connected. To this end, we construct strongly connected graphs by first randomly constructing line graphs as for SG, then randomly adding a few undirected edges to the line graphs, particularly, 30 edges for the 15-robot case and 90 edges for the 45-robot case. At each coordination step, DFS-SG randomly picks the first robot to select an action, and the order of all other robots is determined by a distributed method based on depth-first search. The resulting decision sequence may involve relay robots that transmit information between robots that are not directly connected and, thus, DFS-SG generally requires longer decision times than SG.
Figure 9: Comparative Analysis of Coverage Performance with 45 robots for $100$ Mbps data rate. The reported times include ROS1 delays of up to 3.78 sec per action coordination step.
Figure 10: Comparative Analysis of Coverage Performance with 45 robots for $0.25$ Mbps data rate. The reported times include ROS1 delays of up to 6.54 sec per action coordination step.
@@
  • Network Self-Configuration: This module applies only to RAG since only RAG enables the network to self-configure itself across the planning steps subject to the robots’ bandwidth constraints and the robots’ relative locations. Particularly, as described also in the above paragraph for RAG-$k$nn, at the beginning of each planning step, each robot selects the $k$ nearest other robots within its communication range. This neighborhood selection scheme is justified by Example 1.
  • Perception: The process is different for each algorithm since each algorithm requires the robots to process information received from different sets of robots:
    For RAG, the robots need the following three operations to detect the nearby environment and evaluate where would be best to move: (i) At the beginning of each planning step, each robot hovers at a constant height (30 m) and rotates its camera to take a picture in each of the 8 cardinal directions relative to its body frame. The captured FOV in each direction is of size $34.6\times 26.0$ m2. Then, the robot uses semantic segmentation [^9] We simulate semantic segmentation using the in-built ground truth result of AirSim. The process runs at 20Hz, ensuring segmentation does not create a bottleneck for the replanning frequency of RAG. to detect the road segments in each of the 8 captured images. The resulting segmented image has size 25 KB with label information after compression. (ii) Each robot, upon receiving segmented images and relative poses from neighbors that have committed actions in specific directions, stitches these images together. That way, the robot reconstructs the collective FOV of its neighbors that have committed an action. This reconstruction will be used next by the robot to select in which direction to move to cover the most new road area. To this end, (iii) the robot stitches its 8 captured images respectively with the previous reconstruction and counts the extra road pixels that can be covered in each of the corresponding 8 cardinal directions.
@@
  • Action Coordination: The process differs across algorithms:
    For RAG, given their neighbors’ already committed actions, the robots that are still deciding will first each pick a trajectory whose FOV gives the most amount of newly covered road area. Then, they will compare this amount with one another. Those who win their neighbors will commit to their picked trajectories and share the corresponding FOV and camera pose with neighbors. Otherwise, they will receive the FOVs and camera poses selected by newly committed robots and repeat the process above.
@@ VII-A Evaluations with 15 Robots
We present the results for the 15-robot scenario. Each compared algorithm was tested in 50 trials, each lasting 300 seconds. The simulation setup is as follows.
Simulation Setup. Across the Monte Carlo tests, the robots are initialized near one another such that their FOVs largely overlap (e.g., see the leftmost group of 15 robots in Fig. 1). Therefore, coordination becomes necessary for the robots to spread in the environment such that they detect different road segments and, thus, maximize the total new road area detected after each action coordination step.
@@
Results. The results are summarized in Figs. 7 (100 Mbps) and 8 (0.25 Mbps). RAG-$k$nn achieves real-time planning for all $k$, and superior mean coverage for $k\geq 2$.
Road coverage over time
Figs. 7(a) and 8(a) present the total area of the road covered over time for each algorithm. RAG-$k$nn rapidly achieves superior mean performance for $k\geq 2$, and maintains comparable performance for $k=1$. The lower performance for $k=0$ is expected, since then no robot ever coordinates with any other.
Trade-off between decision time vs. road coverage
Figs. 7(b) and 8(b) present the average action coordination time vs. the peak coverage of the road area over the duration of the task (the peak values in Figs. 7(a) and 8(a)). For larger $k$, the marginal increase in total coverage is negatively outweighed by a greater increase in decision time. The observation demonstrates the trade-off between decision speed vs. optimality predicted in Section V (Remark 5).
VII-B Evaluations with 45 Drones
We demonstrate the scalability of our algorithm from 15 to 45 robots. Each compared algorithm was tested in 30 trials, each lasting 500 seconds. The experiments demonstrate that RAG scales linearly, being at most 3 times slower compared to the 15-robot case. In contrast, the compared near-optimal algorithms scale cubically, as predicted by our theoretical analysis, being $>27=3^{3}$ times slower compared to the 15-robot case: e.g.SG’s decision time increases from around 2 sec to 60 sec (100 Mbps case), a 30x increase, and from around 14 sec to 850 sec (0.25 Mbps case), a 60x increase.
Simulation Setup. Across the Monte Carlo tests, the robots are divided into three groups, with the robots within each group being deployed near to one another such that their FOVs largely overlap (Fig. 1). In such a setting, RAG automatically enables parallelized decision-making, achieving scalability and similar coverage as for the 15-robot setting, in contrast to the state-of-the-art near-optimal algorithms.
@@ VII-B Evaluations with 45 Drones
Results. The results are summarized in Figs. 9 (100 Mbps) and 10 (0.25 Mbps). RAG maintains superior mean coverage, with all qualitative observations from the 15-robot case applying here. Per Figs. 9(b) and 10(b), RAG’s coordination times scaled linearly compared to the 15-robot case. In contrast, SG’s and DFS-SG’s coordination times scaled super-cubically, requiring now minutes per action coordination step at 100 Mbps and tens of minutes at 0.25 Mbps. The results demonstrate the linear scalability of RAG and cubic scalability of the state of the art, as predicted by our theoretical analysis (Table I).
VIII Summary and Concluding Remarks
We introduced a distributed submodular optimization paradigm, Resource-Aware distributed Greedy (RAG), that enables scalable and near-optimal coordination over robot mesh networks. The framework applies to any distributed submodular optimization task. In this paper, we applied it to active information-gathering, demonstrating RAG’s performance in simulated scenarios of road detection with up to 45 robots. In the simulations, RAG enabled real-time planning, up to 3 orders of magnitude faster than competitive near-optimal algorithms, while also achieving superior mean coverage performance. To enable the simulations, we extended the high-fidelity and photo-realistic AirSim by enabling a scalable collaborative autonomy pipeline to tens of robots while simulating realistic r2r communication messages and speeds.
In our future work, we will apply RAG to distributed metric-semantic SLAM where the r2r communication messages scale to MBs [57, 8], 40 times larger than the messages of 25 KB we considered for the road detection task in this paper.
@@ VIII Summary and Concluding Remarks
We also plan the following algorithmic contributions. (i) RAG assumes synchronous communication. Although RAG can be modified to handle such cases, e.g., by instructing each robot to execute its action without first waiting for all other robots to select actions, its near-optimality guarantees provided in this paper may become invalid. Our future work will extend our theoretical and algorithmic analysis beyond the above limitations. (ii) We will also extend our results to handle effective task execution over long time horizons. For example, in collaborative mapping over long time horizons, the team needs to stay updated on the areas that have been mapped, such that the current plans do not repeat past actions. This is in addition to the focus of this paper that only the current plans among the robots do not overlap. To this end, we will handle network connectivity constraints. (iii) Finally, we will enhance our simulator by integrating the simulation of realistic communication channels and protocols towards communication-aware and -efficient coordination algorithms.
@@
$=f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}\cup{\cal B}_{2}})-f(a_{i}\,|\,\{a_{j}\}_{j\,\in\,{\cal B}_{1}})\leq 0,$ (20)
where the inequality holds since $f$ is submodular. Hence, ${\sf coin}_{f,i}({\cal N}_{i})$ is non-increasing in ${\cal N}_{i}$. Therefore, for any $i\in{\cal N}$, ${\sf coin}_{f,i}({\cal N}_{i})$ achieves the lower bound ${\sf coin}_{f,i}({\cal N}_{i})\geq{\sf coin}_{f,i}({\cal N}\setminus\{i\})=f(a_{i})-f(a_{i}\,|\,\emptyset)=0$. For the upper bound,
@@
$=f({\cal A}^{\sf OPT}\cup{\cal A}^{\sf RAG})-\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}^{\sf OPT}\cup{\cal A}_{[i-1]}^{\sf RAG})$ (23)
$\leq f({\cal A}^{\sf RAG})+\sum_{i\in{\cal N}}f(a_{i}^{\sf OPT}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$
$\quad-(1-\kappa_{f})\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$ (24)
$\leq f({\cal A}^{\sf RAG})+\kappa_{f}\sum_{i\in{\cal N}}f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})$ (25)
$=(1+\kappa_{f})f({\cal A}^{{\sf RAG}})$
$\quad+\kappa_{f}\sum_{i\in{\cal N}}\left[f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}\cap[i-1]})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]})\right]$ (26)
$\leq(1+\kappa_{f})f({\cal A}^{{\sf RAG}})+\kappa_{f}\sum_{i\in{\cal N}}\left[f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{[i-1]\setminus\mathcal{N}_{i}})\right]$ (27)
$\leq(1+\kappa_{f})f({\cal A}^{{\sf RAG}})+\kappa_{f}\sum_{i\in{\cal N}}\underbrace{\left[f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG}\,|\,{\cal A}^{{\sf RAG}}_{\mathcal{N}_{i}^{c}})\right]}_{{\sf coin}_{f,i}({\cal N}_{i})},$ (28)
where eq. 23 holds from telescoping the sums; eq. 24 holds from $f$ being submodular and $1-\kappa_{f}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}\setminus\{i\}}^{\sf RAG})}{f(a_{i}^{\sf RAG})}\leq\frac{f(a_{i}^{\sf RAG}\,|\,{\cal A}^{\sf OPT}\cup{\cal A}_{[i-1]}^{\sf RAG})}{f(a_{i}^{\sf RAG}\,|\,{\cal A}_{{\cal N}_{i}\cap[i-1]}^{\sf RAG})}$ per the definition of $\kappa_{f}$; eq. 25 holds since RAG selects $a^{{\sf RAG}}_{i}$ greedily; eq. 26 holds from telescoping the sums; eq. 27 holds from $f$ being 2nd-order submodular; and eq. 28 holds from $f$ being submodular. Therefore, eq. 12 holds.
Then, in the fully centralized case where ${\cal N}_{i}^{c}=\emptyset,\forall i\in{\cal N}$, we have ${\sf coin}_{f,i}({\cal N}_{i})=f(a_{i}^{\sf RAG})-f(a_{i}^{\sf RAG})=0$. Hence, eq. 13 follows eq. 28.
@@
where eq. 29 holds from eq. 8, and eq. 30 holds from [58, Lemma 2.1]. Combining eqs. 28 and 30, the lower bound in eq. 15 can be proved. ∎